OpenStudy (anonymous):

plz help gold medal will be awarded. solve : 6x-2(x+1=3x-10

OpenStudy (anonymous):

Do you mean\[6x - 2(x+1) = 3x - 10\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok so balance the equation, put all x on one side and all numbers on the other. So we can multiply out that bracket to get 6x -2x -2 = 3x -10 Simplify the left hand side 4x -2 = 3x -10 Take away 3x from both sides now x -2 = -10 Add 2 to both sides x = -8 There's your answer :)

OpenStudy (anonymous):

ok r u sure

OpenStudy (anonymous):

Yes 100% that's right!

OpenStudy (anonymous):

ok thx

OpenStudy (anonymous):

No problem :)